# Taming Cross-Domain Representation Variance in Federated Prototype Learning with Heterogeneous Data Domains

### v1

@@ 5 Experiments
5.1 Performance Comparison
We compare our proposed method with the state-of-the-art methods using the Digit-5, Office-10 and DomainNet datasets, as detailed in table 1, table 2 and table 3. Our method demonstrates significant improvements in average accuracy over baseline methods for both datasets. A closer examination of domain-specific results reveals a more pronounced enhancement in performance on domains that are more challenging to learn. For instance, within the Digit-5 dataset, our method achieves a $5.3\%$ increase in accuracy for the SVHN domain, which is more difficult, and a $0.58\%$ increase for the MNIST dataset, considered easier. This aligns with the goal of our method, which is to address the disparate learning challenges across various domains, enhancing fairness and aiding in performance improvement in harder domains. Similar improvements can also be observed in Office-10 and DomainNet.
5.2 Ablation Study
5.2.1 Impact of Dual-Level Prototype Generation.
In this subsection, we focus on examining the impact of dual-level clustered prototypes, as demonstrated by the results in table 4. The first row illustrates outcomes derived from calculating local prototypes by averaging features of samples within the same class, and global prototypes through the direct averaging of these local prototypes. This approach, utilizing straightforward averaging for both levels, yields the least effective performance. This is attributed to its failure to capture variance information, a non-trivial aspect in this context. Moreover, adopting a solely global clustering approach does not significantly enhance performance. This is attributed to the fact that while global clusters incorporate inter-domain variance, they overlook the critical aspect of high intra-domain sample variance, particularly in domains that are challenging to learn. To elucidate this, we present a t-SNE visualization analysis in fig. 3 comparing the three prototype generation methods. Our approach fosters a more generalizable decision boundary, as illustrated in the visualization. This ability to effectively capture variance at both local and global levels is key to why our dual-level clustering method outperforms the others.
Table 4: Comparison on prototype generation methods. Variance means the average distance from the normalized feature vector of one sample to its corresponding class feature center (i.e. the averaged prototype) for each client. Results are then used for visualization in fig. 3. Details in section 5.2.1.
Figure 3: Visualization of different prototype generation methods. The first row averages feature vectors locally and averages local prototypes globally. The second row averages feature vectors locally and clusters local prototypes globally. The last row (ours) clusters feature vectors locally and clusters local clustered prototypes globally. The last column Total is the visualization of mixing the feature vectors from all datasets. Details in section 5.2.1.
To further explore alternative methods, we examine the efficacy of directly transferring all local prototypes to each client without employing any aggregation or clustering techniques. This approach preserves variance information but raises significant privacy concerns, unlike our dual-level clustering method, which offers enhanced privacy protection. As demonstrated in table 5, directly transferring all local prototypes yields performance comparable to our dual-cluster approach. However, it requires transmitting approximately five times as many prototypes in each training round.
Table 5: Comparison between w/o and w/ global clustering. w/o means the server distributes all local clustered prototypes to the clients for local training. Avg #normal-#\## of prototypes is the average number of prototypes each client receives from the server during each global communication round. Details in section 5.2.1.
Figure 4: Impact of $\alpha$ sparsity parameter and $\lambda$ prototype loss weight. The left figure shows the accuracy of two selected datasets (Syth and MNIST-M) and the average accuracy among all clients with different $\alpha$. The right figure shows the effects of different $\lambda$ for both FPL and our proposed approach. Details in section 5.2.3.
Figure 5: Impact of $\tau$. The left figure shows the impact on the Digital-5 dataset while the right figure shows on the Office-10 dataset. Average means the average test accuracy with variance among all clients. Details in section 5.2.2.
5.2.2 Impact of Temperature $\tau$.
Table 6: Comparison on components of $\alpha$-sparsity prototype loss. Contrast and Correction stand for the contrastive and corrective loss term in the total $\alpha$-sparsity loss respectively. Avg is the average accuracy result for all clients. Details in section 5.2.3.
To assess the impact of the contrastive temperature ($\tau$) on our model’s performance, we conduct the following experiment. The results, depicted in fig. 5, demonstrate that our method consistently surpasses the baselines across a range of temperatures (refer to table 1 and table 2 for details, where FPL exhibits the highest baseline performance at $66.88\%$ and $58.55\%$ respectively). Further analysis indicates that an optimal temperature setting for our model on Digit-5 is $\tau=0.070$. Another results on Office-10 identifies an optimal temperature setting for our model at $\tau=0.045$, while we maintain $\tau=0.070$ for stability in performance comparisons. It is noteworthy that temperatures either significantly higher or lower than this value lead to training difficulties due to numerical instability.
5.2.3 Impact of $\alpha$-Sparsity Prototype Loss.
To evaluate the specific impact of our proposed $\alpha$-sparsity prototype loss, we conduct experiments comparing both contrastive and corrective loss terms. Results in table 6 demonstrates that employing the contrast term led to a $1.46\%$ improvement in final average accuracy, while the correction term resulted in a $0.99\%$ enhancement. Combining all components yields the best performance, showcasing a $2.29\%$ improvement. Our investigation also focuses on the sparsity parameter $\alpha$, depicted in fig. 4. We find that varying $\alpha$ within the range of $(0,1)$ consistently outperforms the baseline setting with $\alpha=1$. After careful consideration of numerical stability, we opt for an $\alpha$ value of $0.250$.
6 Conclusion

### v2

@@ Appendix C Further results on DomainNet
Figure 6: Impact of $\alpha$ on the DomainNet dataset.
Figure 7: Test accuracy on DomainNet with different Dirichlet distribution parameters.
Appendix D Label Non-I.I.D. Setting
We expanded our method’s assessment on Digit-5 and Office-10 within a non-i.i.d. label setting. Utilizing the Dirichlet method ($\alpha$ = 0.5), we shaped the data distribution and generated non-i.i.d. datasets for individual clients. Notably, our approach showcased substantial accuracy improvements compared to baseline methods across both datasets. Non-i.i.d. distributions are visualized in Fig. 8, the variation in features across non-identically distributed data is reflected by the colors of the dots, whereas the non-identical label distribution ($\alpha=0.5$) is denoted by the sizes of the dots. For the top figure in Fig. 8, client index $0,1,2,3,4$ own the datasets MNIST, SVHN, USPS, Synth and MNIST-M respectively; for the bottom figure, client index $0,1,2,3$ represent the dataset domains Amazon, Caltech, DSLR and Webcam repectively.
Figure 8: Visualization of data distributions under label non-i.i.d. setting. Every point on the graph symbolizes a collection of samples belonging to a specific class assigned to a client. The left figure is from digit-5 and the right figure is from office-10. Details in Sec. D
The outcomes are presented in Tab. 7 and Tab. 8. Consistent with our findings in the i.i.d. label setting, we observed a significant performance boost in domains that pose greater learning challenges. For instance, within the Digit-5 dataset, our method achieved a $2.24\%$ accuracy surge in the Synth domain—a more complex task—while showing a $0.94\%$ increase in accuracy for the comparatively easier USPS dataset. This reiterates our method’s aim: addressing disparate learning difficulties among domains, fostering fairness, and bolstering performance, especially in tougher domains. Similar advancements were evident within the Office-10 dataset.
Table 7: Test accuracy on Digit-5 under label non-i.i.d. setting. Avg means average results among all clients. We apply the Dirichlet method ($\alpha$ = 0.5) to obtain the data distribution and create the non-i.i.d. dataset for each client. Details in Sec. D.
Methods Digit-5 MNIST SVHN USPS Synth MNIST-M Avg $\Delta$ FedAvg 79.44 $\pm$ 0.86 24.16 $\pm$ 0.91 65.46 $\pm$ 1.86 40.58 $\pm$ 0.87 55.16 $\pm$ 1.24 52.96 - FedProx 81.40 $\pm$ 0.53 24.44 $\pm$ 1.06 67.54 $\pm$ 2.05 41.64 $\pm$ 1.02 55.80 $\pm$ 0.58 54.16 +1.20 FedProto 83.28 $\pm$ 0.80 23.06 $\pm$ 0.51 69.98 $\pm$ 0.81 40.78 $\pm$ 0.99 57.66 $\pm$ 0.85 54.95 +1.99 FedPCL 85.44 $\pm$ 2.13 25.82 $\pm$ 0.97 69.28 $\pm$ 3.73 42.60 $\pm$ 1.55 58.66 $\pm$ 1.46 56.36 +3.40 FedFA 86.65 $\pm$ 0.69 29.57 $\pm$ 1.97 73.48 $\pm$ 0.77 47.19 $\pm$ 0.49 60.24 $\pm$ 1.24 59.43 +6.47 FPL 85.18 $\pm$ 2.66 31.54 $\pm$ 2.96 72.96 $\pm$ 2.41 49.48 $\pm$ 2.42 61.08 $\pm$ 2.10 60.05 +7.09 Ours 86.18 $\pm$ 0.98 33.24 $\pm$ 1.31 73.98 $\pm$ 2.17 51.72 $\pm$ 1.08 62.90 $\pm$ 1.32 61.60 +8.64
Table 8: Test accuracy on Office-10 under label non-i.i.d. setting. Avg means average results among all clients. We apply the Dirichlet method ($\alpha$ = 0.5) to obtain the data distribution and create the non-i.i.d. dataset for each client. Details in Sec. D.
Methods Office-10 Amazon Caltech DSLR Webcam Avg $\Delta$ FedAvg 47.36 $\pm$ 1.60 33.63 $\pm$ 0.21 49.42 $\pm$ 1.37 72.95 $\pm$ 1.38 50.84 - FedProx 48.58 $\pm$ 0.76 33.78 $\pm$ 0.63 52.42 $\pm$ 5.31 75.10 $\pm$ 0.80 52.47 +1.63 FedProto 52.44 $\pm$ 0.86 35.56 $\pm$ 1.09 51.38 $\pm$ 0.19 75.62 $\pm$ 2.11 53.75 +2.91 FedPCL 54.07 $\pm$ 0.76 36.56 $\pm$ 2.18 57.62 $\pm$ 0.11 76.42 $\pm$ 0.82 56.17 +5.33 FedFA 53.28 $\pm$ 1.72 36.67 $\pm$ 1.49 53.71 $\pm$ 1.43 73.20 $\pm$ 2.52 54.22 +3.38 FPL 53.05 $\pm$ 0.50 35.41 $\pm$ 1.05 58.67 $\pm$ 2.95 76.62 $\pm$ 1.60 55.94 +5.10 Ours 55.28 $\pm$ 0.76 37.48 $\pm$ 0.21 59.79 $\pm$ 1.90 77.36 $\pm$ 0.08 57.48 +6.64
Appendix E Different Clustering Algorithms
In our paper, FINCH is employed for our dual-level prototype clustering due to its parameter-free nature. Here we select the one with the minimum number of cluster centers among several possible clustering schemes generated by FINCH. This characteristic obviates the need for hyper-parameter tuning typically required in clustering methods, such as selecting the number of centers in K-Means algorithm. However, this does not imply exclusivity to the FINCH clustering algorithm within our method. Our approach is compatible with various clustering algorithms, including simpler methods like K-Means. The experiments result in Tab. 9 show that with a carefully tuned K-Means algorithm (precisely determining the appropriate number of centers), our method with K-Means achieves performance comparable to that achieved using FINCH. Conversely, poorly tuned parameters result in weaker performance with the K-Means clustering. This reinforces our choice of FINCH, for its parameter-free advantage.
Table 9: Comparison with K-Means Algorithm. Adaptive K means we use the number of clustering centers from FINCH as K. Details in Sec. E.
Local Global MNIST SVHN USPS Synth MNIST-M Avg Adaptive K Adaptive K 89.97 42.10 85.83 60.90 66.03 68.97 K = 2 K = 2 89.53 40.67 85.33 59.13 63.53 67.64 K = 5 K = 5 89.72 41.97 85.41 60.13 64.69 68.38 FINCH FINCH 90.70 42.08 86.24 60.08 67.16 69.25
Appendix F Unbalanced Clients Distribution Setting
We further explore the performance of our FedPLVM in handling unbalanced client distributions, mirroring the scenario demonstrated in FPL. Within a pool of 10 clients, the data ownership is distributed as follows: 1 client possesses MNIST domain data, 4 clients hold SVHN domain data, 2 clients possess USPS domain data, 2 clients hold Synth domain data, and 1 client owns MNIST-M domain data.
The outcomes, presented in Tab. 10, showcase our approach’s equitable performance, notably in enhancing test accuracy on clients dealing with the more challenging datasets. Comparatively, in contrast to FPL, our method displays a significant $2.94\%$ improvement in SVHN dataset performance. It is worth noting that our ‘avg’ column in Tab. 10 represents the average test accuracy across all clients, differing from FPL, which calculates the average test accuracy first among clients with the same dataset and then averages across datasets. This distinction addresses potential unfairness, wherein a higher representation of clients with ‘easy’ datasets could disproportionately influence the final accuracy in FPL.
Table 10: Comparison on unbalanced clients distribution. Test accuracy on each dataset domain is the average result among all clients that own the corresponding dataset. Avg means average results among all clients. Details in Sec. F.
Methods MNIST SVHN USPS Synth MNIST-M Avg FPL 90.44 $\pm$ 0.67 59.78 $\pm$ 1.56 85.14 $\pm$ 0.93 73.01 $\pm$ 1.09 69.34 $\pm$ 0.98 71.52 Ours 90.90 $\pm$ 0.97 62.72 $\pm$ 1.69 85.92 $\pm$ 0.75 74.51 $\pm$ 1.20 71.32 $\pm$ 0.35 73.40
Appendix G Privacy Protection
In our method, the server only receives clustered local prototypes, making it challenging to reconstruct the clients’ local datasets. We also employ a differential privacy (DP) technology to validate the impact of FedPLVM. Each client trains local model by DP-SGD [1] to perturb model parameters. The noise multiplier is determined by [22, 4, 3]. The privacy budget $\epsilon$ and approximate parameter $\delta$ are set as $4.0$ and $51e-$ 5 respectively for a $(\epsilon,\delta)$-DP setting. Meanwhile, we also incorporate the clustered local prototypes with the privacy protection technique. We set the scale parameter $s=0.05$ and the perturbation coefficient $p=0.1$ for the Gaussian noise distribution generation of our local clustered prototypes. As shown in Tab. 11, the average accuracy drops at most $0.77\%$ if we employ the privacy protection only for the prototypes. The approximate DP causes a slightly larger decrease in accuracy, up to $2.45\%$. Note that even with the privacy protection technologies, our method reaches a comparable performance to the most advanced baseline in Tab. 1.
Table 11: Impact of differential privacy. Avg means average results among all clients. w/ and w/o represents we incorporate the local model or the local clustered prototypes with the privacy protection technologies or not. Details in Sec. G.
Model Prototypes MNIST SVHN USPS Synth MNIST-M Avg w/ w/ 89.74 $\pm$ 0.35 38.40 $\pm$ 2.86 84.44 $\pm$ 0.48 55.98 $\pm$ 1.27 62.44 $\pm$ 1.25 66.20 w/ w/o 89.42 $\pm$ 0.85 38.12 $\pm$ 1.12 83.68 $\pm$ 0.43 57.54 $\pm$ 1.74 65.26 $\pm$ 0.51 66.80 w/o w/ 90.64 $\pm$ 0.34 40.12 $\pm$ 1.95 86.40 $\pm$ 0.99 58.80 $\pm$ 1.53 66.44 $\pm$ 1.60 68.48 w/o w/o 90.70 $\pm$ 0.39 42.08 $\pm$ 1.59 86.24 $\pm$ 1.37 60.08 $\pm$ 1.47 67.16 $\pm$ 0.77 69.25
Appendix H Global Prototypes Generation
The determination of the number of global prototypes for a class is contingent on the associated learning challenges. In instances where a class within a domain presents significant learning difficulties, it is characterized by large variance. Consequently, employing a single global prototype, reducing the sufficiency of variance information, could decay the learning performance. Incorporating multiple global prototypes for such a class can introduce additional variance information, thereby enhancing the learning process. We provide averaged experiment results in Tab. 12. (Note in the main paper we only report the single experiment to be consistent with the corresponding visualization.) The results validate the necessity of our dual-level clustering.
Table 12: Comparison with Local Cluster & Global Average. Details in Sec. H.
Local Global MNIST SVHN USPS Synth MNIST-M Avg
Cluster Avg 89.43 $\pm$ 0.55 40.57 $\pm$ 1.21 85.74 $\pm$ 0.89 59.77 $\pm$ 0.88 65.03 $\pm$ 1.02 68.11
Cluster Cluster 90.70 $\pm$ 0.39 42.08 $\pm$ 1.59 86.24 $\pm$ 1.37 60.08 $\pm$ 1.47 67.16 $\pm$ 0.77 69.25
Appendix I Local Prototypes Clustering
Local prototype clustering is beneficial because it captures essential variance information, not just the average data, which is particularly crucial in hard domains. Easy domains tend to show tight clustering within the same category and clear distinctions between different categories, facilitating accurate classification. In contrast, hard domains often exhibit looser clustering, increasing the risk of misclassification, especially for samples near category boundaries. Therefore, only capturing average data suffices for easy domains but falls short for hard domains.
Our proposed method captures more feature distribution variation information by introducing the local prototypes clustering, since clustering provides more prototypes compared to simply averaging especially considering the sparse distribution in hard domains as shown in Fig. 3. For easy domains, where the average is sufficient, our method generates fewer prototypes.
To demonstrate this, we refer to Fig. 9. The y-axis shows the average number of local prototypes among classes generated at each selected round. The easy domain (MNIST) has fewer prototypes generated compared to the hard domain (SVHN), showing that in hard domains, more prototypes are utilized to better capture the variance information. Furthermore, an average performance gain of $3.08\%$ in experiment of impact of local prototypes clustering in Tab. 3 also supports this observation.
Figure 9: Tendency of average number of local clustered prototypes for different classes in different domains. Details in Sec. I.
Appendix J Discussion on Wider Range of Scenarios
J.1 Limitation.
While FedPLVM introduces innovative strategies to address domain heterogeneity in federated learning, it also comes with certain limitations and challenges. For example, noisy labels, or incorrectly labeled data instances, are a common issue in real-world datasets. A few previous works has discussed the related impact in FL scenario [37, 38, 36]. This noisy case will bring two problems for prototype learning in FL: Prototype Distortion. Noisy labels can lead to distorted prototype representations, as the local prototypes clustered for each class may be influenced by mislabeled instances. This distortion can propagate throughout the training process, affecting the model’s ability to accurately capture the true underlying data distribution. Prototype Ambiguity. In the presence of noisy labels, the distinction between different classes becomes less clear, leading to ambiguity in prototype definitions. Prototypes may no longer accurately represent the true characteristics of their respective classes, making it difficult for the model to generalize effectively. These make noisy labels in FL scenario an interesting direction for future extension to our proposed FedPLVM.
J.2 Broader Impact.
FedPLVM primarily addresses the challenges of cross-domain federated learning, where data from heterogeneous domains need to be collaboratively utilized while preserving privacy. This capability enables the application of FedPLVM in various real-life scenarios, including: Healthcare Data Collaboration. Different hospitals or healthcare institutions often possess diverse patient datasets with varying characteristics and distributions. FedPLVM allows these institutions to collaboratively train machine learning models for tasks such as disease diagnosis, patient outcome prediction, or personalized treatment recommendation while ensuring data privacy and security. Internet of Things (IoT) Networks. IoT devices deployed in different locations or environments generate heterogeneous data streams, including sensor readings, environmental data, and user interactions. FedPLVM enables federated learning across IoT networks, facilitating tasks such as anomaly detection, predictive maintenance, or environmental monitoring without centralized data collection, thus preserving user privacy and reducing communication overhead.